OpenStudy (anonymous):

I will give a medal! :) Will someone please help me solve this? Solve -4(x + 1) - 3 = -3(x - 4). -19 19 2 0

OpenStudy (lxelle):

-4x-4-3=-3x+12 -4x+3x=12+4+3 -x=19 x=19

OpenStudy (anonymous):

let me try and work this out for myself thank you! :)

OpenStudy (lxelle):

Sure. :)

OpenStudy (anonymous):

I am confused can you please explain in steps?

OpenStudy (unklerhaukus):

expand the brackets \[ -4(x + 1) \longrightarrow -4\times x-4\times1\]

OpenStudy (anonymous):

thank you! I think I can work the rest out for myself